Please help me solve this equation :
(x - 2)² -x^2+8=0

Answer by Cromlix(4381)

Hi there,
(x - 2)² -x^2+8=0
Multiply out bracket.
(x^2 - 4x + 4) - x^2 + 8 = 0
Remove brackets.
x^2 - 4x + 4 - x^2 + 8 = 0
The x^2 cancel out
-4x + 4 + 8 = 0
-4x + 12 = 0
Move the 12 over to the other side
remembering to change the sign.
-4x = -12
Divide both sides by -4
x = -12/-4
x = 3.
Hope this helps :-)